# Momentum–mass normalized dark–bright solitons to one dimensional Gross–Pitaevskii systems

### v2

@@ 4 Properties of the minimizing surface
In particular, $(1-\tilde{\rho}^{2})^{2}\in L^{2}(\mathbb{R})$. In addition, Theorem 3.5 implies
@@
is well-defined, $\gamma\in(0,1]$ and $p(\tilde{u})=q$. Furthermore, Proposition 3.2 assures that $v^{\star}\in L^{2}(\mathbb{R})$ with $\|v^{\star}\|_{2}^{2}=m$. Hence, $(\tilde{u},v^{\star})\in\mathcal{X}_{q,m}$.
@@
Letting $\varepsilon$ tend to zero, we deduce that the function $q\mapsto E_{\textup{min}}(q,m)$ is nondecreasing in $0,\infty)$. Furthermore, the fact that $E_{\textup{min}}(0,m)=0$ for all $m\geq 0$ implies that $E_{\textup{min}}$ is nonnegative. This finishes the proof of (i).
@@
Case 2: Assume now that $q_{1}=q_{2}$. Since (2.8) is trivially satisfied for $q_{1}=q_{2}=0$, we may assume in addition that $q_{1}=q_{2}>0$. Arguing as for the previous case, let $(u_{1},v_{1})\in\mathcal{X}_{q_{1},m_{1}}^{0}$ be such that $E(u_{1},v_{1})<E_{\textup{min}}(q_{1},m_{1})+\varepsilon$. In addition, let $v\in\mathcal{C}_{0}^{\infty}(\mathbb{R})$ be such that $\|v\|_{2}^{2}=m_{2}-m_{1}$ and $E(1,v)<\varepsilon$ (one may find such a function $v$ by rescaling a given function with mass $m_{2}-m_{1}$, as in the proof of Lemma 4.3). Applying a translation if necessary, assume that $v$ and $v_{1}$ have disjoint supports, and define $v_{2}=v_{1}+v$. Then, $\|v_{2}\|_{2}^{2}=m_{2}$ and
@@
The following result provides the boundedness of almost minimizing sequences.
@@
Next result establishes a suitable positive bound from below for minimizing sequences. This is a key step for proving a subadditivity property for $E_{\textup{min}}$. The second part of the lemma will also be used to rule out vanishing in the concentration-compactness argument, see Remark 5.3.
@@
$E(u_{n},v_{n})\geq E(u_{n},0)-\frac{\alpha}{2}\int_{\mathbb{R}}(1-\rho_{n}^{2})v_{n}^{2}dx\geq E_{\textup{min}}(q,0)-\frac{\alpha}{2}\int_{\mathbb{R}}(1-\rho_{n}^{2})v_{n}^{2}dx.$
@@
For $j=1,2$, let $\{(u_{n}^{(j)},v_{n}^{(j)}\}\subset\mathcal{X}_{q_{j},m_{j}}^{0}$ be such that
@@
Recall that $0\leq 1-\rho_{n}^{(j)}<1$ for $j=1,2$. Since they have disjoints supports, it follows that $0\leq 1-\rho_{n}^{(1)}+1-\rho_{n}^{(2)}<1$ too. Proposition 3.2 implies that $\|1-\rho_{n}^{(1)}+1-\rho_{n}^{(2)}\|_{\infty}=\|(1-\rho_{n}^{(1)}+1-\rho_{n}^{(2)})^{\star}\|_{\infty}$, so $\rho_{n}>0$. Taking now $\theta_{n}$ as any primitive of $\zeta_{n}$, we define $u_{n}=\rho_{n}e^{i\theta_{n}}$.
Observe that (2.10) trivially holds for $q_{1}=q_{2}=0$, being both sides equal to zero. Hence, from now on we assume that $q_{1}+q_{2}>0$. In such a case, taking into account that the supports are disjoint, from Proposition 3.3 and Theorem 3.4 we deduce
@@
$k_{n}=\frac{3}{4}\Big{(}\min\{\|(\rho_{n}^{(1)})^{\prime}\|_{2},\|(\rho_{n}^{(2)})^{\prime}\|_{2}\}+\min\{\|(v_{n}^{(1)})^{\prime}\|_{2},\|(v_{n}^{(2)})^{\prime}\|_{2}\}\Big{)}.$ (4.9)
@@
Therefore, recalling (4.9), we deduce that $k_{n}>3\delta^{2}/(4\sqrt{2}(m_{1}+m_{2}))$ for all $n$. This concludes the proof. ∎
@@ 5 Compactness of minimizing sequences
Next lemma provides some weak and local convergences of almost minimizing sequences. It is a preliminary step towards the strong convergence of $v_{n}$ and $1-|u_{n}|^{2}$ in $L^{2}(\mathbb{R})$.
@@
Next theorem leads to the compactness of almost minimizing sequences and, in turn, to the existence of a constrained energy minimizer.
@@
Let $\alpha>0$, $\beta\geq 0$, $q\in(0,\pi/2)$, and $m>0$. Assume that either (H1) or (H2) holds. Let $\{(u_{n},v_{n})\}\subset\mathcal{X}$ satisfy (4.5), as well as
@@
We emphasize that, in the proof of Theorem 5.2, the fact that $m>0$ allows to apply Lemma 4.5, which is the key to rule out vanishing. Naturally, the result is also true for $m=0$ (with $v\equiv 0$), even though the proof of the compactness of almost minimizing sequences in this case differs from ours, see [4].
@@
We also note that the compactness of almost minimizing sequences is often a powerful tool in establishing orbital stability, following the approach of [10]. Although this method is not directly applicable in our setting, since the fixed quantity $p(u)$ is not conserved along the flow in general, we have chosen to state and prove the result in the broader context of almost minimizing sequences, rather than restricting ourselves to truly minimizing sequences, in order to emphasize that this step does not present an obstacle to proving stability.
@@
Thus, $E(u,v)=E_{\textup{min}}(q,m)$ provided that $(u,v)\in\mathcal{X}_{q,m}$. We thus claim that $(u,v)\in\mathcal{X}_{q,m}$, i.e. $p(u)=q$ and $\|v\|_{2}^{2}=m$. The proof follows by a concentration-compactness argument as in [2]. Indeed, for every $n$, let us denote $u_{n}=\rho_{n}e^{i\theta_{n}}$ and consider the function
@@
In order to estimate the right-hand side of (5.17), using (5.15), one has
@@
$\int_{\mathbb{R}}$ $\Big{|}(\rho_{n}^{\prime})^{2}-((\rho_{n}^{(1)})^{\prime})^{2}-((\rho_{n}^{(2)})^{\prime})^{2}\Big{|}\,\mathrm{d}x\leq C\int_{\{t_{n}<|x|<8t_{n}\}}\Big{(}(\rho_{n}^{\prime})^{2}+(1-\rho_{n})^{2}\Big{)}\,\mathrm{d}x$
@@
Then, [34, Lemma 1.2] implies that
@@
Therefore, applying [34, Lemma 3.3], we conclude that that $v_{n}\to 0$ in $L^{4}(\mathbb{R})$. This contradicts Lemma 4.5.
@@
Then, given $\varepsilon>0$, a standard procedure provides some $t_{\varepsilon}>0$ such that
@@ 6 Proof of the main result
$\begin{cases}\displaystyle-\rho^{\prime\prime}+\frac{c^{2}(1-\rho^{4})}{4\rho^{3}}=(1-\rho^{2}-\alpha v^{2})\rho,&x\in\mathbb{R},\\ -v^{\prime\prime}=(\lambda-\alpha\rho^{2}-\beta v^{2})v,&x\in\mathbb{R}.\end{cases}$ (6.1)
Arguing by contradiction, assume that there exists $x_{0}\in\mathbb{R}$ such that $v(x_{0})=0$. Since $v\geq 0$ in $\mathbb{R}$ and $v\in\mathcal{C}^{1}(\mathbb{R})$, it follows that $v^{\prime}(x_{0})=0$. Thus, applying Cauchy-Lipschitz theorem to the second equation in (6.1) implies that $v\equiv 0$, a contradiction. Therefore, $v>0$ in $\mathbb{R}$.

### v3

@@ 4 Properties of the minimizing surface
In particular, $1-\tilde{\rho}^{2}\in L^{2}(\mathbb{R})$. In addition, Theorem 3.5 implies
@@
is well–defined, $\gamma\in(0,1]$ and $p(\tilde{u})=q$. Furthermore, Proposition 3.2 assures that $v^{\star}\in L^{2}(\mathbb{R})$ with $\|v^{\star}\|_{2}^{2}=m$. Hence, $(\tilde{u},v^{\star})\in\mathcal{X}_{q,m}$.
@@
Letting $\varepsilon$ tend to zero, we deduce that the function $q\mapsto E_{\textup{min}}(q,m)$ is nondecreasing in $0,\infty)$.
On the other hand, given $(u,v)\in\mathcal{E}(\mathbb{R})\times H^{1}(\mathbb{R})$, Young inequality and $\alpha^{2}\leq\beta$ imply
$\frac{1}{4}(1-|u|^{2})^{2}+\frac{\beta}{4}v^{4}-\frac{\alpha}{2}(1-|u|^{2})v^{2}\geq\frac{1}{4}\Big(1-\frac{\alpha^{2}}{\beta}\Big)(1-|u|^{2})^{2}\geq 0.$
Hence, $E(u,v)\geq 0$, so that $E_{\textup{min}}(q,m)\geq 0$ for all $q\in[0,\pi/2)$, $m\geq 0$.
In addition, let $m\geq 0$ and $v\in H^{1}(\mathbb{R})$ such that $\|v\|_{2}^{2}=m$. For every $\lambda>0$, define $v_{\lambda}(x)=\sqrt{\lambda}v(\lambda x)$, for all $x\in\mathbb{R}$. It is clear that $\|v_{\lambda}\|_{2}^{2}=m$ for all $\lambda>0$, and
$E(1,v_{\lambda})=\int_{\mathbb{R}}\left(\frac{\lambda^{2}}{2}(v^{\prime})^{2}+\frac{\beta\lambda}{4}v^{4}\right)\,\mathrm{d}x.$
Therefore $E(1,v_{\lambda})\to 0$ as $\lambda\to 0$. In particular, since $(1,v_{\lambda})\in\mathcal{X}_{0,m}$ for all $\lambda>0$, it follows that $E_{\textup{min}}(0,m)=0$. This finishes the proof of (i).
@@
Case 2: Assume now that $q_{1}=q_{2}$. Since (2.8) is trivially satisfied for $q_{1}=q_{2}=0$, we may assume in addition that $q_{1}=q_{2}>0$. Arguing as for the previous case, let $(u_{1},v_{1})\in\mathcal{X}_{q_{1},m_{1}}^{0}$ be such that $E(u_{1},v_{1})<E_{\textup{min}}(q_{1},m_{1})+\varepsilon$. In addition, let $v\in\mathcal{C}_{0}^{\infty}(\mathbb{R})$ be such that $\|v\|_{2}^{2}=m_{2}-m_{1}$ and $E(1,v)<\varepsilon$ (one may find such a function $v$ by rescaling a given function with mass $m_{2}-m_{1}$, as in the proof of item (i) or the proof of Lemma 4.3). Applying a translation if necessary, assume that $v$ and $v_{1}$ have disjoint supports, and define $v_{2}=v_{1}+v$. Then, $\|v_{2}\|_{2}^{2}=m_{2}$ and
@@
The following result provides the boundedness of sequences whose energy converges to $E_{\textup{min}}(q,m)$ and whose mass converges to $m$.
@@
Next result establishes a suitable positive bound from below for minimizing sequences. This is a key step for proving a subadditivity property for $E_{\textup{min}}$. The second part of the lemma will also be used to rule out vanishing in the concentration–compactness argument, see Remark 5.4.
@@
$E(u_{n},v_{n})\geq E(u_{n},0)-\frac{\alpha}{2}\int_{\mathbb{R}}(1-\rho_{n}^{2})v_{n}^{2}{\,\mathrm{d}x}\geq E_{\textup{min}}(q,0)-\frac{\alpha}{2}\int_{\mathbb{R}}(1-\rho_{n}^{2})v_{n}^{2}{\,\mathrm{d}x}.$
@@
If $q_{1}q_{2}=0$, then (2.10) is a direct consequence of item (iii) in Theorem 2.3. Thus, let us consider the case $q_{1}q_{2}>0$. For $j=1,2$, let $\{(u_{n}^{(j)},v_{n}^{(j)}\}\subset\mathcal{X}_{q_{j},m_{j}}^{0}$ be such that
@@
Recall that $0\leq 1-\rho_{n}^{(j)}<1$ for $j=1,2$. Since they have disjoints supports, it follows that $0\leq 1-\rho_{n}^{(1)}+1-\rho_{n}^{(2)}<1$ too. Proposition 3.2 implies that $\|1-\rho_{n}^{(1)}+1-\rho_{n}^{(2)}\|_{\infty}=\|(1-\rho_{n}^{(1)}+1-\rho_{n}^{(2)})^{\star}\|_{\infty}$, so $\rho_{n}>0$. Taking now $\theta_{n}$ as any primitive of $\zeta_{n}$, we define $u_{n}=\rho_{n}e^{i\theta_{n}}$. Taking into account that the supports are disjoint, from Proposition 3.3 and Theorem 3.4 we deduce
@@
$k_{n}={\frac{3}{8}}\Big(\min\{\|(\rho_{n}^{(1)})^{\prime}\|_{2},\|(\rho_{n}^{(2)})^{\prime}\|_{2}\}+\min\{\|(v_{n}^{(1)})^{\prime}\|_{2},\|(v_{n}^{(2)})^{\prime}\|_{2}\}\Big).$ (4.9)
@@
Therefore, recalling (4.9), we deduce that $k_{n}>3\delta^{2}/({8}\sqrt{2}(m_{1}+m_{2}))$ for all $n$. This concludes the proof. ∎
@@ 5 Compactness of minimizing sequences
Next lemma provides some weak and local convergences of sequences satisfying (4.5). It is a preliminary step towards the strong convergence of $v_{n}$ and $1-|u_{n}|^{2}$ in $L^{2}(\mathbb{R})$.
@@
Remark 5.2.
We underline that hypotheses (H1) and (H2) are only required to prove (5.6), and then (5.7), in Lemma 5.1. In some previous works on the Gross–Pitaevskii equation with nonzero conditions at infinity, the lack of lower bounds of type (5.6) has been overcome by introducing different notions of the momentum which take values in $\mathbb{R}\setminus\pi\mathbb{Z}$, and then fixing this new momentum in the minimization argument instead of our renormalized momentum (2.1), see [5, 11, 12, 33]. Although allowing the dark component to vanish could potentially yield a more general result without the assumptions (H1) and (H2), it is not clear how to adapt an approach such as ours, based on symmetric decreasing rearrangements, if liftings of the dark components are not allowed and, in turn, the energy cannot be written in terms of $\rho$ and $\theta^{\prime}$.
@@
Next theorem leads to the compactness of sequences satisfying (4.5) and, in turn, to the existence of a constrained energy minimizer.
@@
Let $\alpha>0$ and $\beta>0$ satisfy $\alpha^{2}\leq\beta$ and let $q\in(0,\pi/2)$ and $m>0$. Assume that either (H1) or (H2) holds. Let $\{(u_{n},v_{n})\}\subset\mathcal{X}$ satisfy (4.5), as well as
@@
We emphasize that, in the proof of Theorem 5.3, the fact that $m>0$ allows to apply Lemma 4.5, which is the key to rule out vanishing. Naturally, the result is also true for $m=0$ (with $v\equiv 0$), even though the proof of the compactness in this case differs from ours, see [4].
@@
We also note that the compactness of sequences satisfying (4.5)–(5.11) is often a powerful tool in establishing orbital stability, following the approach of [10]. Although this method is not directly applicable in our setting, since the fixed quantity $p(u)$ is not conserved along the flow in general, we have chosen to state and prove the result in the broader context of sequences satisfying (4.5)–(5.11), rather than restricting ourselves to truly minimizing sequences, in order to emphasize that this step does not present an obstacle to proving stability.
Remark 5.6.
We point out that, in the proof of Theorem 5.3, the condition $\alpha^{2}\leq\beta$ is used only to have the strict inequality in (2.10).
@@
Thus, $E(u,v)=E_{\textup{min}}(q,m)$ provided that $(u,v)\in\mathcal{X}_{q,m}$. We thus claim that $(u,v)\in\mathcal{X}_{q,m}$, i.e. $p(u)=q$ and $\|v\|_{2}^{2}=m$. The proof follows by a concentration–compactness argument as in [2], which in turn is based on the classical method in [28] (see also [9]). Indeed, for every $n$, let us denote $u_{n}=\rho_{n}e^{i\theta_{n}}$ and consider the function
@@
In order to estimate the right–hand side of (5.17), using (5.15), one has
@@
$\int_{\mathbb{R}}$ $\Big|(\rho_{n}^{\prime})^{2}-((\rho_{n}^{(1)})^{\prime})^{2}-((\rho_{n}^{(2)})^{\prime})^{2}\Big|\,\mathrm{d}x\leq C\int_{\{t_{n}<|x|<8t_{n}\}}\Big((\rho_{n}^{\prime})^{2}+{(1-\rho_{n}^{2})^{2}}\Big)\,\mathrm{d}x$
@@
Then, [37, Lemma 1.2] implies that
@@
Therefore, applying [37, Lemma 3.3], we conclude that $v_{n}\to 0$ in $L^{4}(\mathbb{R})$. This contradicts Lemma 4.5.
@@
Equivalently, $\mu_{0}=\mu$. Then, given $\varepsilon>0$, a standard procedure (see for instance Step 1 of the proof of [9, Proposition 1.7.6]) provides some $t_{\varepsilon}>0$ such that
@@ 6 Proof of the main result
$\begin{cases}\displaystyle-\rho^{\prime\prime}+\frac{c^{2}(1-\rho^{4})}{4\rho^{3}}=(1-\rho^{2}-\alpha v^{2})\rho,&x\in\mathbb{R},\\ -v^{\prime\prime}=(\lambda-\alpha\rho^{2}-\beta v^{2})v,&x\in\mathbb{R},\end{cases}$ (6.1)
and moreover, $2\theta^{\prime}=c(\rho^{-2}-1),$ where $u=\rho e^{i\theta}$. Arguing by contradiction, assume that there exists $x_{0}\in\mathbb{R}$ such that $v(x_{0})=0$. Since $v\geq 0$ in $\mathbb{R}$ and $v\in\mathcal{C}^{1}(\mathbb{R})$, it follows that $v^{\prime}(x_{0})=0$. Thus, applying Cauchy–Lipschitz theorem to the second equation in (6.1) implies that $v\equiv 0$, a contradiction. Therefore, $v>0$ in $\mathbb{R}$.
@@
Appendix
In this last section we prove that the bright components are always real–valued modulo a constant change of phase.
Proposition 6.1.
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuous function, and assume that there exists $\lambda\in\mathbb{R}$ such that $f(x)\leq\lambda$ for all $x\in\mathbb{R}$. Let $v\in H^{1}(\mathbb{R};\mathbb{C})$ be a solution to
$v^{\prime\prime}=f(x)v,\,x\in\mathbb{R}.$ (6.5)
Then, there exists $\theta_{0}\in\mathbb{R}$ such that $v(x)e^{i\theta_{0}}\in\mathbb{R}$ for all $x\in\mathbb{R}$.
Remark 6.2.
Given any solution $(u,v)\in\mathcal{E}(\mathbb{R})\times H^{1}(\mathbb{R};\mathbb{C})$ to the Gross–Pitaevskii system, one may apply Proposition 6.1 with $f(x)=\lambda-\alpha|u(x)|^{2}-\beta|v(x)|^{2}$ and derive that $v$ is real–valued up to a multiplication by a constant of modulus one.
Proof of Proposition 6.1.
Let us assume that $v$ is not identically zero, since the result is trivial otherwise. Let $(a,b)$ be a possibly unbounded interval such that
$|v(x)|>0,\,\text{for all }x\in(a,b),\quad\lim_{x\to a^{+}}|v(x)|=\lim_{x\to b^{-}}|v(x)|=0.$
In such an interval, we may take polar coordinates $v(x)=\rho(x)e^{i\theta(x)}$, where $\rho$ and $\theta$ are $\mathcal{C}^{2}$ in $(a,b)$. Notice that $\theta$ may not have limits at $a$ or $b$. Writing the equation (6.5) in terms of $\rho$ and $\theta$, and equating real and imaginary parts, we derive
$-\rho^{\prime\prime}+(\theta^{\prime})^{2}\rho=f(x)\rho,\quad$ $x\in(a,b),$ (6.6)
$2\rho^{\prime}\theta^{\prime}+\rho\theta^{\prime\prime}=0,\quad$ $x\in(a,b).$ (6.7)
Equation (6.7) is equivalent to $(\rho^{2}\theta^{\prime})^{\prime}=0$ in $(a,b)$, which implies that there exists $C\in\mathbb{R}$ such that
$\rho^{2}\theta^{\prime}=C,\quad x\in(a,b).$ (6.8)
We claim that $C=0$. To prove the claim, let us substitute (6.8) into (6.6) to obtain
$\rho^{\prime\prime}+\frac{C^{2}}{\rho^{3}}=f(x)\rho,\quad x\in(a,b).$ (6.9)
In what follows we distinguish between the cases of $(a,b)=\mathbb{R}$ or $(a,b)\not$$\mathbb{R}$.
Let us first assume that $(a,b)=\mathbb{R}$. Using that $f\leq\lambda$ in $\mathbb{R}$, and rearranging terms in (6.9), we obtain
$\rho^{\prime\prime}\geq\frac{C^{2}}{\rho^{3}}-\lambda\rho,\quad x\in\mathbb{R}.$ (6.10)
Seeking a contradiction, assume that $C^{2}>0$. Let $\varepsilon>0$ be such that
$\frac{C^{2}}{\varepsilon^{3}}-\lambda\varepsilon:=R_{\varepsilon}>0.$
Let $x_{\varepsilon}\in\mathbb{R}$ be such that $\rho(x)<\varepsilon$ for all $x\in[x_{\varepsilon},\infty)$. Then, from (6.9) we deduce
$\rho^{\prime\prime}(x)\geq R_{\varepsilon},\quad x\in[x_{\varepsilon},\infty).$ (6.11)
Then, integrating twice, we conclude
$\rho(x)\geq\frac{R_{\varepsilon}}{2}(x-x_{\varepsilon})^{2}+\rho^{\prime}(x_{\varepsilon})(x-x_{\varepsilon})+\rho(x_{\varepsilon}),\quad x\in[x_{\varepsilon},\infty).$
We arrive at a contradiction by letting $x$ tend to infinity. Hence, $C=0$ and, from (6.8), it follows that there exists $\theta_{0}\in\mathbb{R}$ such that $\theta=-\theta_{0}$ in $\mathbb{R}$, i.e. $v(x)e^{i\theta_{0}}=|v(x)|\in\mathbb{R}$ for all $x\in\mathbb{R}$. Thus, the result is proved in the case $(a,b)=\mathbb{R}$.
We consider now the case $(a,b)\not$$\mathbb{R}$. Assume first that $a\in\mathbb{R}$. Then, $v$ satisfies
$\begin{cases}-v^{\prime\prime}=f(x)v,&x\in(a,b),\\ v(a)=0.\end{cases}$ (6.12)
If we denote $v=v_{1}+iv_{2}$, it is clear that $v_{j}$ satisfies (6.12) too for $j=1,2$. Since the second–order equation in (6.12) is linear, it is known that the vector space of real–valued solutions to the equation has dimension 2, and since $v_{1}(a)=v_{2}(a)=0$, there exists $\mu\in\mathbb{R}$ such that $v_{1}=\mu v_{2}$ in $(a,b)$. This implies that $|v|=\sqrt{1+\mu^{2}}|v_{2}|$ in $(a,b)$. Notice that $v_{2}$ cannot change sign in $(a,b)$ since, otherwise, $|v|$ would vanish at some point in $(a,b)$. Therefore, we may assume that $v_{2}>0$ in $(a,b)$, an thus,
$\rho=|v|=\sqrt{1+\mu^{2}}v_{2},\quad x\in(a,b).$
Multiplying equation (6.9) by $(1+\mu^{2})^{-1/2}$, it follows
$v_{2}^{\prime\prime}+\frac{C^{2}}{(1+\mu^{2})^{2}v_{2}^{3}}=f(x)v_{2},\quad x\in(a,b).$
Bearing in mind that $v_{2}$ satisfies also $v_{2}^{\prime\prime}=f(x)v_{2}$ in $(a,b)$, we conclude again that $C=0$. Therefore, there exists $\theta_{0}\in\mathbb{R}$ such that $v(x)e^{i\theta_{0}}\in\mathbb{R}$ for all $x\in(a,b)$.
It is clear that the function $w:=e^{i\theta_{0}}v$ is a solution to (6.5), it is real–valued and does not vanish in $(a,b)$, and $w(a)=0$. If $w^{\prime}(a)=0$, the uniqueness of solution to the Cauchy problem associated to $w^{\prime\prime}=f(x)w$ would imply that $w\equiv 0$, a contradiction. As a consequence, $w^{\prime}(a)\not=0$. This implies that $w$ is real–valued also in $(a-\varepsilon,a)$ for some $\varepsilon>0$. Then, arguing as above, it can be proved that $w$ is real–valued in the possibly unbounded interval $(a_{1},a)$ such that $|w|>0$ in $(a_{1},a)$ and $\lim_{x\to a_{1}^{+}}|w(x)|=0$. An argument by induction proves that $w$ is real–valued in $(-\infty,b)$. If $b=\infty$, the proof is finished, while if $b\in\mathbb{R}$, the proof follows the same lines as we showed for $a\in\mathbb{R}$. ∎